I am trying to help my daughter understand how to solve a word problem but can't remember how to solve it myself -
"If sales of 30 cattle bring in a profit of $4500 and sales of 50 cattle bring in a profit of $12,500, and the profit model is linear, how much profit will the sale of 75 cattle bring in?"
$4,500 divided by 30 = $150 per head
$12,500 divided by 50 = $250 per head
my thought was that linear would mean 75 x $350 = $26,250 but that was not one of the multiple choice answers

Found 3 solutions by stanbon, scott8148, solver91311:
Answer by stanbon(75887)

If sales of 30 cattle bring in a profit of $4500 and sales of 50 cattle bring in a profit of $12,500, and the profit model is linear, how much profit will the sale of 75 cattle bring in?
---
You have two points relating # of cattle and profit:
(30,4500) and (50,12500)
---
The linear equation has the form p = mc + b
where p is profit and c is number of cattle.
-------------------
m is the slope of the equation:
m = (12500-4500)/(50-30) = 8000/20 = 400
-----
Find "b" using p=4500 when c=30 and m = 400:
4500 = 400*30 + b
b = 4500-12000
b = -7500
----
Equation:
p(c) = 400c - 7500
----
Solve: "how much profit will the sale of 75 cattle bring in?"
p(75) = 400*75 - 7500
p(75) = $22,500
==================
Cheers,
Stan H.

Answer by scott8148(6628)

the linear profit model means that the income per steer (slope) is constant , but there is also overhead or fixed costs (intercept)

20 more cattle (50 - 30) bring in 8000 more dollars (12500 - 4500)
so the income per steer is ___ 8000/20 or 400

30 cattle means income of 12000 , but only 4500 profit
this indicates a fixed cost (overhead) of 7500 ___ this checks out for 50 cattle

the equation is ___ p = 400c - 7500

plug in 75 for c to find your answer

Answer by solver91311(24713)



This is nothing more than creating an equation given two ordered pairs and then using the equation to predict the location of a third point. Instead of the normal x-y coordinate plane, you have a set of axes labeled 'Profit' on the vertical and 'Cattle sold' on the horizontal. A point on your plane would then be identified by an ordered pair (c, p) where c is the number of cattle sold and p is the corresponding profit.

You are given two points, namely (30,4500) and (50,12500)

To derive your equation, use the two-point form of an equation of a straight line, namely:



where and are the coordinates of the given points.

Just fill in the numbers and do the necessary arithmetic to simplify your equation. Once you have the equation written, substitute 75 for x and calculate the value of y, that is the profit that corresponds. Let me know what you get for an answer.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism